OpenStudy (abbycross167):

Can someone please help me fix this math question? I will give a medal and will fan... Question -

OpenStudy (abbycross167):

Factor by the trinomial method - x^2 – 8x + 7 My steps - Step 1: Find a, b, and c A = 1 B = -8 C = 7 Step 2: Use b and c to determine what two numbers will go at the end of the two parentheses. (x -/+ ?)(x -/+ ?) Find factors of c (7) that add up to b (-8). Since 7 can be written as the product of 7 and 1, and since 7+1=8, then use 7 and 1. Step 3: Answer format: (variable _ #)(variable _ #) Write in the two numbers found in Step 2 above: (x+7)(x+1) Answer: x2 – 8x + 7 = (x+7)(x+1)

OpenStudy (abbycross167):

What my teacher said was wrong - "But the 8 is negative!"

OpenStudy (abbycross167):

@bmk614 do you think you could try to help me fix this quesion?

OpenStudy (zenmo):

Well, what does (x+7)(x+1) give you?

OpenStudy (zenmo):

You can check your answer by using the FOIL method.

OpenStudy (zenmo):

\[x^2+8x+7\], notice that the 2nd term is positive where the original equation 2nd term is negative -- hence, your answer is wrong.

OpenStudy (abbycross167):

oh okay... would you mind helping me to fix it?

OpenStudy (zenmo):

(x-7)(x-1)

OpenStudy (abbycross167):

Is that what the answer needs to be?

OpenStudy (zenmo):